Suppose you construct a simple linear regression model. Using 5 observations the LSE of the slope is calculated as 0.3 with associated standard error .86. Now you wish to test H0:β=0 versus H1:β≠0 at the 0.05 significance level. Applying an appropriate test, you

reject null

fail to reject null

either reject or fail to reject null

have no conclusion
Answer #1

here as test statsitc =estimated slope/standard error =0.3/0.86 =0.349

as at 0.05 level and (n-2=3) degree of freedom ; critical t =3.182

for test statistic is not in rejection region

fail to reject null